5. At a given time a shaft is rotating 50 rpm about a fixed axis, 20 s later it is rotating 1050 rpm. Solve its angular acceleration.
1
Expert's answer
2020-07-10T10:22:08-0400

By definition, the angular acceleration is:


"\varepsilon = \dfrac{\Delta \omega}{\Delta t}"

where "\Delta \omega" is angular velocity change during the time "\Delta t".

The "\Delta \omega" can be calculated as following (as far as one revolution corresponds to "2\pi" radians):

"\Delta \omega = (1050rpm - 50rpm)\cdot 2\pi = 2000\pi \space rad/m \approx 104.7\space rad/s"

Thus, the angular acceleration will be:


"\varepsilon = \dfrac{104.7\space rad/s}{20s} \approx 5.2 rad/s^2"

Answer. 5.2 rad/s^2.